# Evaluating the Impact of Inter-cluster Communications in Edge Computing

### v1

@@ Abstract
Distributed applications based on micro-services in edge computing are becoming increasingly popular due to the rapid evolution of mobile networks. While Kubernetes is the default framework when it comes to orchestrating and managing micro-service-based applications in mobile networks, the requirement to run applications between multiple sites at cloud and edge poses new challenges. Since Kubernetes does not natively provide tools to abstract inter-cluster communications at the application level, inter-cluster communication in edge computing is becoming increasingly critical to the application performance. In this paper, we evaluate for the first time the impact of inter-cluster communication on edge computing performance by using three prominent, open source inter-cluster communication projects and tools, i.e., Submariner, ClusterLink and Skupper. We develop a fully open-source testbed that integrates these tools in a modular fashion, and experimentally benchmark sample applications, including the ML class of applications, on their performance running in the multi-cluster edge computing system under varying networking conditions. We experimentally analyze two classes of envisioned mobile applications, i.e., a) industrial automation, b) vehicle decision drive assist. Our results show that Submariner performs best out of the three tools in scenarios with small payloads, regardless of the underlying networking conditions or transmission direction between clusters. When sending larger data to a service, ClusterLink outperforms Submariner once the inter-node networking conditions deteriorate, which may be the case in highly mobile scenarios in edge computing. Finally, Skupper significantly outperforms others in a variety of scenarios with larger payloads.
@@ I Introduction
The inter-cluster communication in the edge context is in fact in its infancy. In this context, research to date majorly focuses on either the placement of workloads across the infrastructure[1], orchestration of multiple clusters[2] or the methods for cluster creation based on various properties, such as node proximity[3]. However, research on how these components can horizontally scale by actually establishing connectivity between them and how this impacts application performance is lacking. Static IP connectivity, while feasible for cloud environments, would require a significant maintenance overhead in dynamic environments like MEC. In fact, such an approach is not suited for large-scale and evolving deployments, therefore necessitating flexible connectivity creation based on service name resolution across cluster boundaries. As we face a rapidly emerging space of open-source projects that address the area of inter-cluster communication, the question arises as to how to assess and evaluate their impact on the application performance in edge environments.
In this work, we experimentally study and comparatively analyze for the first time three notable open source projects and tools for inter-cluster communication in edge networking and computing: Submariner, ClusterLink and Skupper. While there are several other tools for general inter-cluster communication, this study focuses on open-source solution that do not necessitate sidecar containers, or a specific container network interface in the connected clusters, and as such allow for the highest degree of compatibility with existing infrastructure. To this end, we engineer an open-source testbed for inter-cluster communication tools. Its realization uses only production-grade open-source tools, while it provides the ability to vary the underlying network condition to mimic an edge network infrastructure setting with various link qualities. The inter-cluster communication tools are implemented in a pluggable fashion, thus allowing for a comparative analysis. We also design a set of innovative case studies for novel distributed mobile applications, including ML-based, which necessitate inter-cluster communication. Finally, we provide measurements and evaluate the impact on application performance. In the measurements, we attempt to answer the following three research questions, which remain unaddressed by existing work for the scope of inter-cluster communication: 1. How does the distance between clusters impact the service response time for latency sensitive applications? 2. What is the combined effect of cluster distance and payload size on the application’s performance? 3. How do various inter-cluster communication tools comparatively perform for latency sensitive applications?
@@ II-A Related work
Kubernetes is the de facto default tool for managing workloads in the cloud. It organizes computing resources in so-called clusters that can be formed based on different characteristics of nodes, administrative or management goals, or yet a combination of them. Dynamic clusters are formed based on aspects subject to continuous changes, like node proximity in scenarios with moving computing nodes [1]. Static clusters are usually based on hardware properties to form specialized clusters, for example, to process tasks that demand hardware accelerators such as GPUs [4]. Specialized clusters are employed to guarantee the compatibility of applications that demand specific resources since those clusters might leverage enhancements to utilize specific resources like GPUs[5]. Additionally, scheduling and workload sharing can differ in those particular clusters, which could lead to incompatibility with other Kubernetes clusters [6]. General clustering approaches are analyzed in [3], showing how dependent the clustering objectives can be on the specific use case.
@@
Distributed application and service function chains can be formed across multiple clusters. The applications are scheduled and scaled based on node mobility or network properties to minimize service latency[2, 7], increase cost-efficiency, reduce power consumption or boost robustness [3]. Multi-cluster deployments of applications are gaining attention because they address a variety of issues faced in distributed environments. On the one hand, there are business reasons to rely on multiple cloud providers (i.e., , multi-cluster environments), such as preventing vendor lock-in and ensuring a global presence. On the other hand, there are technical reasons, such as the need for edge computing to fulfill stringent latency requirements of novel applications (e.g., extended reality (ER)) [8], creating scalability issues. In  [9], the authors show that the Kubernetes control plane can become a bottleneck when a certain cluster size is exceeded. This restriction limits the size of clusters since consensus has to be maintained across all member nodes. While the issue can be reduced by changing or improving the used key-value storage system, this does not provide a complete solution, and the problem persists with a higher geographical spread of the nodes. Additionally, due to the inherent characteristics of edge nodes (i.e., constrained computing and storage capacity), some applications can only partially be executed in an edge cluster, which in turn requires communication with other clusters.
@@
In general, research focusing on the inter-cluster communication problem to allow multi-cluster collaboration is work in progress. In [10], the authors perform an assessment of Submariner’s performance as well as the distributed application deployed on top of it with a focus on CPU utilization. However, this study lacks consideration of two critical dimensions: varying inter-node/inter-cluster latency and the amount of data being exchanged between the clusters. In [11] the authors explore inter-cluster communication by proposing a framework to deploy a 5G experimental network across distributed infrastructure formed by multiple clusters spread geographically. The authors evaluated the maximum achievable throughput in inter-cluster communications provisioned by Submariner using three different tunneling options (i.e., Wireguard, Libreswan, and VXLAN) and three different transport protocols (TCP, UDP and SCTP). These works focus on Submariner while our focus is on a comparative analysis in various scenarios with ClusterLink and Skupper.
@@
Focusing on cloud-driven scenarios for mobile networks furthermore, paper [12] proposed a solution for multi-cluster management and connectivity employing two open-source tools: KubeFed and NSM. VXLAN technology was used to tunnel traffic between clusters. The approach proposed utilized TCP, UDP, and SCTP as transport protocols and benchmarked basic data plane performance, collecting maximum throughput and latency metrics between containers. Since latency between nodes was static, real-world performance in the network edge would diverge significantly from these findings. A valuable extension of paper [12] would in fact be to compare the tool proposed with the tools that we use in our paper, which are production grade.
@@
Related work in benchmarking of intra-cluster networking solutions, such as in [13] focus on total performance within a single cloud cluster, which varies based on the hardware used. Paper [14] conducted an extensive survey about benchmarking performance at the edge and observed that further research is needed to measure the performance of edge networks efficiently. This survey reveals a gap in explicitly benchmarking orchestrators and other software that connects the infrastructure into an edge continuum. Indeed, many benchmarks are limited by physical aspects of edge devices (  e.g., processing power, chipset throughput, storage, and memory). Our evaluation differs from previous works as it focuses on the relative response time differences encountered through varying aspects of 1) underlying network, 2) payload size, and 3) direction of payload exchange between clusters.
@@ II-B Inter-cluster Communication Tools
Tab. I lists inter-cluster communication tools that we implement in this paper; Submariner [^1] https://submariner.io/ (CNCF, RedHat) and Skupper [^2] https://skupper.io/, along with the emerging ClusterLink [^3] https://clusterlink.net/ (IBM) [15]. While all three of them tackle the same challenge that is of inter-cluster service communication, each tool has unique features that set it apart. Skupper, being a layer 7 interconnect, does not need any further configuration of the CIDR ranges that are used to assign IP addresses to pods or services, making it scale very well with comparably lower configuration effort. Submariner on the other hand employs a broker for connection establishment, leading to only this one component having to be publicly accessible; it can potentially even be hosted on hardware that is not part of the communicating clusters. This allows for configurationless cluster-to-cluster communication through NAT as well as firewalls. ClusterLink in turn provides authentication as well as granular permission and access management on workload, service and cluster level, allowing for a high degree of security and low levels of implicit trust between endpoints. All three solutions are written in Go, provide state-of-the-art encryption out of the box and are deployed on top of the respective Kubernetes clusters without the need for modifications of the individual node’s operating systems.
TABLE I: Comparison of multiple cluster connectivity tools
Tool name Deployed Network Features
@@
Skupper.io 1.7.3 TLS Pod IP range agnostic
Submariner 0.18.0 Libreswan/VXLAN/ NAT compatible
Wireguard
ClusterLink v0.4.0 mTLS ACLs
@@ II-C1 Remote Access, Monitoring, and Maintenance
In an industrial automation setting, numerous sensors installed on production machines distributed across a factory continuously collect data about the status of machines in the production line. A massive amount of raw data is cached in a cluster of edge nodes for preprocessing, while aggregating the data where possible. The result of local processing is sent to another cluster for evaluation (e.g., based on machine learning), with the result potentially being used to instruct a control center to perform the required maintenance [16, 17].
@@ II-C2 Vehicle Decision Assist
A specific vehicle that is part of a fleet of vehicles platooning on the road decides to overtake the vehicle in front, for example to optimize fuel consumption, avoid breaking when a preceding vehicle leaves the platoon at an upcoming exit, or because the vehicle in front is facing a mechanical problem. Due to road and traffic conditions, the vehicle cannot overtake easily, so the platoon asks for information to assist with the maneuver. The exchange of information happens between the cluster formed by vehicles in the platoon and a cluster of Road Side Units or oncoming vehicles of the opposite direction (e.g., another platoon). The vehicles share speed, length, and position information to guarantee a safe overtaking. Since reliability and safety is of utmost importance, the overall system consumes information from multiple systems. In this use case, multi-cluster communication is required since the platooning vehicles and the roadside units supporting the vehicles would be grouped in different clusters [18, 19].
Table II illustrates how we extract measurable KPIs for the applications chosen, including references to the literature that we base these KPIs on.
@@
and Maintenance[20, 16, 21]
Vehicle Decision Assist[18] 1000 100
@@ III-A Operating System and Kubernetes
To provide a container orchestration layer, we choose the k3s Kubernetes distribution meant to be used for lightweight and edge devices, v1.30.3+k3s1. It provides the majority of features available in the regular Kubernetes version while having a much smaller footprint. Two clusters are formed with one worker and one headnode each and a total of four different IP ranges between the two clusters for their respective services and pods. As can be seen in Fig. 1, the worker nodes run the application and load generator respectively, while the headnodes are used exclusively to run each cluster’s control plane. This separation ensures that the impact of control plane tasks on the workers is minimized, while also replicating real-world use cases where at least one node has to be available via a defined address to allow for management and negotiation of connection endpoints.
@@ III-C Applications
This function accepts any payload and echoes the same message back to the sender by setting the response body to the received payload. It is used in testing our Remote maintenance inference use case, since it emulates the transfer of aggregated data from one cluster to another. In the test, the result is transmitted back to the sending cluster, which is the behavior that this function copies.
@@
This function receives any payload sent as the HTTP body and acknowledges it with an HTTP 200 response but no reply body. It is used for the Remote maintenance training application, since we can see the same data transfer here as in the previous function with the difference that the data is not transferred back. This behavior would be expected whenever an ML model is trained, since the sending cluster would not have to wait for the training to complete.
@@ IV-A1 Remote maintenance inference
In this class of applications, the payload-echo function emulates the transmission of aggregated data to a specialized cluster for processing. The destination cluster communicates back to the aggregation cluster. Therefore, data has to be transmitted in both directions, where, depending on the specific data and actuators, the amount per direction could fluctuate significantly. For simplification, the destination cluster echoes back the received data, which should approximate the system behavior described in the use case. In Fig. 2, the payload size of the request and the response are therefore identical. During the execution of this test, the KPI for the response time is considered to be 1000 ms. Once this cannot be ensured by the different tools, potentially catastrophic failure in a production system can be assumed.
@@
IV-A2 Remote maintenance ML model training
@@ IV-A3 Vehicle decision assist
As in all tests, we evaluate the results to determine with which inter-cluster network parameter set the response times break the maximum value to ensure save operation. Here, we set this value to 100 ms before the information is deemed so outdated that it loses its meaning and endangers safe operation of the machines.
@@ V Measurements and Results
where $T_{reqSend}$ is the request sending time including possible retransmissions, $T_{hostProc}$ is the host processing time at the destination cluster which includes the time taken by (de-)fragmentation and reassembly of the request. $T_{appProc}$ is the delay introduced by the processing of the application including potential processing of the payload, which depends on the parameter size in bytes. In practice, the differences introduced by different parameter sizes in our experiments are compensated for by the baseline measurement in our results. Finally, $T_{replySend}$ is the reply sending time including possible delays introduced by retransmission mechanisms in the network layer.

### v2

@@ Abstract
Distributed applications based on micro-services in edge computing are becoming increasingly popular due to the rapid evolution of mobile networks. While Kubernetes is the default framework when it comes to orchestrating and managing micro-service-based applications in mobile networks, the requirement to run applications between multiple sites at cloud and edge poses new challenges. Since Kubernetes does not natively provide tools to abstract inter-cluster communications at the application level, inter-cluster communication in edge computing is becoming increasingly critical to the application performance. In this paper, we evaluate for the first time the impact of inter-cluster communication on edge computing performance by using three prominent, open source inter-cluster communication projects and tools, i.e., Submariner, ClusterLink and Skupper. We develop a fully open-source testbed that integrates these tools in a modular fashion, and experimentally benchmark sample applications, including the ML class of applications, on their performance running in the multi-cluster edge computing system under varying networking conditions. We experimentally analyze two classes of envisioned mobile applications, i.e., a) industrial automation, b) vehicle decision drive assist. Our results show that ClusterLink performs best out of the three tools in scenarios with increased payloads, regardless of the underlying networking conditions or transmission direction between clusters. It is closely followed by Skupper, unless request and reply both transport significant amounts of data. Finally, when requesting smaller amounts of data from a service, Submariner slightly outperforms Skupper and ClusterLink regardless of the inter-node networking conditions.
@@ I Introduction
The inter-cluster communication in the edge context is in fact in its infancy. In this area, research to date majorly focuses on either the placement of workloads across the infrastructure[1], orchestration of multiple clusters[2] or the methods for cluster creation based on various properties, such as node proximity[3]. However, research on how these components can horizontally scale by actually establishing connectivity between them and how this impacts application performance is lacking. Static IP connectivity, while feasible for cloud environments, would require a significant maintenance overhead in dynamic environments like MEC. In fact, such an approach is not suited for large-scale and evolving deployments, therefore necessitating flexible connectivity creation based on service name resolution across cluster boundaries. As we face a rapidly emerging space of open-source projects that address the area of inter-cluster communication, the question arises as to how to assess and evaluate their impact on the application performance in edge environments.
In this work, we experimentally study and comparatively analyze for the first time three notable open source projects and tools for inter-cluster communication in edge networking and computing: Submariner, ClusterLink and Skupper. While there are several other tools for general inter-cluster communication, this study focuses on open-source solution that do not necessitate sidecar containers or a specific container network interface in the connected clusters, and as such allow for the highest degree of compatibility with existing infrastructure. To this end, we engineer an open-source testbed for inter-cluster communication tools. Its realization uses only production-grade open-source tools, while it provides the ability to vary the underlying network conditions to mimic an edge network infrastructure setting with various link qualities. The inter-cluster communication tools are implemented in a pluggable fashion, thus allowing for a comparative analysis. We also design a set of innovative case studies for novel distributed mobile applications, including use cases based on machine learning (ML), which necessitate inter-cluster communication. Finally, we provide measurements and evaluate the impact on application performance. In the measurements, we attempt to answer the following three research questions, which remain unaddressed by existing work for the scope of inter-cluster communication: 1. How does the distance between clusters impact the service response time of latency sensitive applications? 2. What is the combined effect of cluster distance and payload size on the application’s performance? 3. How do various inter-cluster communication tools comparatively perform for latency sensitive applications?
@@ II-A Related work
Kubernetes is the de facto default tool for managing workloads in the cloud. It organizes computing resources in so-called clusters that can be formed based on different characteristics of nodes, administrative or management goals, or yet a combination of them. Dynamic clusters are formed based on aspects subject to continuous changes, like node proximity in scenarios with moving computing nodes [1]. Static clusters are usually based on hardware properties to form specialized groups, for example, to process tasks that demand hardware accelerators such as GPUs [4]. Such specialized clusters are employed to guarantee the compatibility of applications that demand specific resources since those devices might leverage enhancements to utilize specific resources like GPUs[5]. Additionally, scheduling and workload sharing can differ in those particular installations, which could lead to incompatibility with other Kubernetes clusters [6]. General clustering approaches are analyzed in [3], showing how dependent the selection objectives can be on the specific use case.
@@
Distributed application and service function chains can be formed across multiple clusters. The applications are scheduled and scaled based on node mobility or network properties to minimize service latency[2, 7], increase cost-efficiency, reduce power consumption or boost robustness [3]. Multi-cluster deployments of applications are gaining attention because they address a variety of issues faced in distributed single-cluster environments. On the one hand, there are business reasons to rely on multiple cloud providers (i.e., , multi-cluster environments), such as preventing vendor lock-in and ensuring a global presence. On the other hand, there are technical reasons, such as the need for edge computing to fulfill stringent latency requirements of novel applications (e.g., extended reality) [8], creating scalability issues. In  [9], the authors show that the Kubernetes control plane can become a bottleneck when a certain cluster size is exceeded. This restriction limits the size of clusters since consensus has to be maintained across all member nodes. While the issue can be reduced by changing or improving the used key-value storage system, this does not provide a complete solution, and the problem can be expect to increase relative to the geographical spread of the nodes. Additionally, due to the inherent characteristics of edge nodes (i.e., constrained computing and storage capacity), some applications can only partially be executed in an edge cluster, which in turn requires communication with other clusters.
@@
In general, research focusing on the inter-cluster communication problem to allow multi-cluster collaboration is work in progress. In [10], the authors perform an assessment of Submariner’s performance as well as the distributed application deployed on top of it with a focus on CPU utilization. However, this study lacks consideration of two critical dimensions: varying inter-node/inter-cluster latency and the amount of data being exchanged between the clusters. In [11] the authors explore inter-cluster communication by proposing a framework to deploy a 5G experimental network across distributed infrastructure formed by multiple geographically spread clusters. The authors evaluated the maximum achievable throughput in inter-cluster communications provided by Submariner using three different tunneling options (i.e., Wireguard, Libreswan, and VXLAN) and three different transport protocols (TCP, UDP and SCTP). These works focus on Submariner while our focus is on a comparative analysis in various scenarios with ClusterLink and Skupper.
@@
Focusing on cloud-driven scenarios for mobile networks furthermore, paper [12] proposed a solution for multi-cluster management and connectivity employing two open-source tools: KubeFed and NSM. The Libreswan technology was used to tunnel traffic between clusters. The approach proposed utilized TCP, UDP, and SCTP as transport protocols and benchmarked basic data plane performance, collecting maximum throughput and latency metrics between containers. Since latency between nodes was static, real-world performance in the network edge would diverge significantly from these findings. A valuable extension of paper [12] would in fact be to compare the tool proposed with some of the production-grade tools that we use in our paper.
@@
Related work in benchmarking of intra-cluster networking solutions, such as in [13] focuses on total performance within a single cloud cluster, which varies based on the hardware used. Paper [14] conducted an extensive survey of benchmarking performance at the edge and observed that further research is needed to measure the performance of edge networks efficiently. This survey reveals a gap in explicitly benchmarking orchestrators and other software that connects the infrastructure into an edge continuum. Indeed, many benchmarks are limited by physical aspects of edge devices (e.g., processing power, chipset throughput, storage, and memory). Our evaluation differs from previous works as, instead of absolute metrics like resource utilization, it focuses on the relative response time differences encountered through varying aspects of 1) underlying network, 2) payload size, and 3) direction of payload exchange between clusters.
@@ II-B Inter-cluster Communication Tools
Tab. I lists inter-cluster communication tools that we implement in this paper; Submariner [^1] https://submariner.io/ (CNCF, RedHat) and Skupper [^2] https://skupper.io/, along with the emerging ClusterLink [^3] https://clusterlink.net/ (IBM) [15]. While all three of them tackle the same challenge that is of inter-cluster service communication, each tool has unique features that set it apart. Skupper, being a layer 7 interconnect, does not need any further configuration of the CIDR ranges that are used to assign IP addresses to pods or services, making it scale very well with comparably low configuration effort. Submariner on the other hand works on layer 3 and employs a broker for connection establishment, leading to only this one component having to be publicly accessible; it can potentially even be hosted on hardware that is not part of the communicating clusters. This allows for cluster-to-cluster communication between dynamic endpoints through NAT as well as firewalls. Using HTTP, ClusterLink shares the benefits of layer 7 connections but also provides authentication as well as granular permission and access management on workload, service and cluster level, allowing for a high degree of security and low levels of implicit trust between endpoints. All three solutions are written in Go, provide state-of-the-art encryption out of the box and are deployed on top of the respective Kubernetes clusters without the need for modifications of the individual node’s operating systems.
TABLE I: Comparison of cluster connectivity tools
Tool name Deployed Network Protocol Features
@@
Skupper.io 1.7.3 mTLS AMQP simple configuration
Submariner 0.18.0 Libreswan IPsec NAT compatible
VXLAN VXLAN
Wireguard Wireguard
ClusterLink v0.4.0 mTLS HTTP ACLs
@@ II-C1 Remote Access, Monitoring, and Maintenance
In an industrial automation setting, numerous sensors installed on production machines distributed across a factory continuously collect data about the status of machines in the production line. A massive amount of raw data is cached in a cluster of edge nodes for preprocessing that aggregate the data where possible. The result of the local processing is sent to another cluster for evaluation (e.g., using machine learning), with the result potentially being used to instruct a control center to perform the required maintenance [16, 17, 20, 21].
@@ II-C2 Vehicle Decision Assist
A specific vehicle that is part of a fleet of vehicles platooning on the road decides to overtake the vehicle in front, for example to optimize fuel consumption, avoid breaking when a preceding vehicle leaves the platoon at an upcoming exit, or because the vehicle in front is facing a mechanical problem. Due to road and traffic conditions, the vehicle cannot overtake easily, so the platoon asks for external information to assist with the maneuver. The exchange of information happens between the cluster formed by vehicles in the platoon and a cluster of Road Side Units or oncoming vehicles of the opposite direction (e.g., another platoon). The vehicles share speed, length, and position information to guarantee a safe overtaking. Since reliability and safety is of utmost importance, the overall system consumes information from multiple systems. In this use case, multi-cluster communication is required since the platooning vehicles per direction and the roadside units supporting the vehicles would all be grouped in different clusters [18, 19].
Table II illustrates the extracted measurable KPIs for the applications chosen, including references to the literature that we base these KPIs on.
@@
and Maintenance[16, 17, 20, 21]
Vehicle Decision Assist[18, 19] 1000 100
@@ III-A Operating System and Kubernetes
To provide a container orchestration layer, we choose the k3s Kubernetes distribution meant to be used for lightweight and edge devices, v1.30.3+k3s1. It provides the majority of features available in the regular Kubernetes version while having a much smaller footprint. Two clusters are formed with one worker and one headnode each and a total of four different IP ranges between the two clusters for their respective services and pods. As can be seen in Fig. 1, the worker nodes run the application and load generator respectively, while the headnodes are used exclusively to run each cluster’s control plane. This separation ensures that the impact of control plane tasks on the workers is minimized, while also replicating real-world use cases where at least one node per cluster has to be available via a defined address to allow for management and negotiation of connection endpoints.
@@ III-C Applications
This function accepts any payload and echoes the same message back to the sender by setting the response body to that received payload. It is used in testing our Remote Maintenance (ML Inference) use case, since it emulates the transfer of aggregated data from one cluster to another with an equally sized response. In the test, the result shall be transmitted back to the sending cluster, which is the behavior that this function exhibits.
@@
This function receives any payload sent as the HTTP body and acknowledges it with an HTTP 200 response without body. It is used for the Remote Maintenance (ML Training) application, since we can see the same data transfer here as in the previous function with the difference that the data is not transferred back. This behavior would be expected whenever an ML model is trained, since the sending cluster would not have to wait for the training to complete.
@@ IV-A1 Remote maintenance inference
In this class of applications, the payload-echo function emulates the transmission of aggregated data to a specialized cluster for processing. The destination cluster communicates back to the aggregation cluster. Therefore, data has to be transmitted in both directions, where, depending on the specific data and actuators, the amount per direction could fluctuate significantly. For simplification, our destination cluster echoes back the received data, which should approximate the system behavior described in the use case. In Fig. 2, the payload size of the request and the response are therefore identical. During the execution of this test, the KPI for the response time is considered to be 1000 ms. Once this cannot be ensured by the different tools, potentially catastrophic failure in a production system can be assumed.
@@
IV-A2 Remote Maintenance (ML Training)
@@ IV-A3 Vehicle decision assist
As in all tests, we evaluate the results to determine with which inter-cluster network parameter set the response times break the maximum value to ensure save operation. Here, we set this value to 100 ms before the information is deemed so outdated that it loses its meaning and endangers safe operation of the machines and safe transportation of the passengers.
@@ V Measurements and Results
where $T_{reqSend}$ is the request sending time including possible retransmissions, $T_{hostProc}$ is the host processing time at the destination cluster which includes the time taken by (de-)fragmentation and reassembly of the request. $T_{appProc}$ is the delay introduced by the processing of the application including potential processing of the payload, which depends on the parameter size in bytes. In practice, the differences introduced by the processing of different parameter sizes in our experiments are compensated for by the baseline measurements (loc0 scenario) in our results. Finally, $T_{replySend}$ is the reply sending time including possible delays introduced by retransmission mechanisms in the network layer.